OpenStudy (anonymous):

Solve by using the substitution method. 6m + n = 34 and m - 5n = 47 Type as an ordered pair.

OpenStudy (zarkon):

what are x,y?

OpenStudy (saifoo.khan):

lol, TYPO! >.<

OpenStudy (saifoo.khan):

m = 7, n = -8

OpenStudy (zarkon):

ok ;)

OpenStudy (anonymous):

n = 34 - 6m m - 5*(34-6m)=47 m = 7 n = 34 - 6*7 = -8 ans: (7, -8)

myininaya (myininaya):

how do you know if is it (m,n) or (n,m)

OpenStudy (zarkon):

good question

OpenStudy (saifoo.khan):

Great question.

OpenStudy (saifoo.khan):

loll, we are giving medals for the GOOD QUESTION.

OpenStudy (anonymous):

I am trying....stupid connection!!!